Determine the energy of 1.10 mol of photons for each of the following kinds of light. (Assume three significant figures.)

part a. visible light (490 nm )

part b. ultraviolet radiation (150 nm )
Answer #1

a)

Given:

lambda = 4.9*10^-7 m

Find energy of 1 photon first

use:

E = h*c/lambda

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(4.9*10^-7 m)

= 4.057*10^-19 J

This is energy of 1 photon

Energy of 1 mol = energy of 1 photon * Avogadro's number

= 4.057*10^-19*6.022*10^23 J/mol

= 2.443*10^5 J/mol

= 2.443*10^2 KJ/mol

Energy of n mol = n * energy of one mol

Energy of 1.1 mol = 1.1* 2.443*10^2

Energy of 1.1 mol = 2.687*10^2 KJ

Answer: 267 KJ

b)

Given:

lambda = 1.5*10^-7 m

Find energy of 1 photon first

use:

E = h*c/lambda

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(1.5*10^-7 m)

= 1.325*10^-18 J

This is energy of 1 photon

Energy of 1 mol = energy of 1 photon * Avogadro's number

= 1.325*10^-18*6.022*10^23 J/mol

= 7.98*10^5 J/mol

= 7.98*10^2 KJ/mol

Energy of n mol = n * energy of one mol

Energy of 1.1 mol = 1.1* 7.98*10^2

Energy of 1.1 mol = 8.778*10^2 KJ

Answer: 878 KJ